how many kilograms of water must be added to 12kg of a 30% salt solution to produce a 20% solution?
Answer by alicealc(293)

water means 0% salt solution
let water = x kilograms

0%*x + 30%*12 = 20%*(x + 12)
(omit all the % sign)
0*x + 30*12 = 20*(x + 12)
0 + 360 = 20x + 240
360 - 240 = 20x
120 = 20x
x = 120/20 = 6

so we must add 6 kilograms of water